One integer is eleven less than two times another. If the sum of the integers is 13, find the two integers.

Answer by sdmmadam@yahoo.com(530)

One integer is eleven less than two times another. If the sum of the integers is 13, find the two integers.
Let the two integers be a and b respectively.
One integer is 11 less than 2 times another.
That is a is 11 less than 2b
which implies a = 2b-11 -----(1)
The sum of the integers in 13
That is a + b = 13 ----(2)
Putting (1) in (2)
(2b-11) + b = 13
(2b + b) = 13 + 11 (change side then change sign)
3b = 24
b = 8
And since a + b = 13 with b = 8
as 5 added to 8 is 13 we have a = 5
Answer: The integers are 5 and 8
Verification:
We should have the first integer a, eleven less than twice the second integer b
Therefore 2b-11 = 2 X 8 - 11 = 16 - 11 = 5 = a which is correct